In a town election, the winning candidate had 220 more votes than the loser. If 810 votes were cast in all, how many votes did each candidate receive?
Answer by checkley71(8403)

X+(X+220)=810 OR 2X=810-220 OR 2X=590 OR X=590/2 OR X=295 VOTES FOR THE LOOSER
810-295=515 VOTES FOR THE WINNER.